# Memory Storyboard: Leveraging Temporal Segmentation for Streaming Self-Supervised Learning from Egocentric Videos

### v1

@@ 5.1 Experiment Setup
We use two real-world egocentric video datasets in the experiments: (1) the child S subset of SAYCam dataset [55], which contains 221 hours of video data collected from a head-mounted camera on the child from age 6-32 months, decoded at 25 fps; (2) the KrishnaCam dataset [52], which contains 70 hours of video data spanning nine months of the life of a graduate student, decoded at 10 fps. There two datasets have also been adopted in a number of existing self-supervised learning literature [40, 43, 75, 59].
@@
Following the architectural choices of Osiris [73], we use ResNet-50 [27] as the feature extractor with group normalization [64] and the Mish activation function [35]. Unless otherwise specified, the default hyperparameter values we use in our experiments are $b=64$, $B=512$, $T=4.5K$ for SAYCam and $T=1.8K$ for KrishnaCam (both corresponding to 3 minutes of raw video), subsampling rate $r=8$ for SAYCam and $r=4$ for KrishnaCam. We train the models with two sets of memory sizes to evaluate their performance across different memory constraints: a larger memory constraint with $|M|=50K$, $|M_{short}|=5K$, $|M_{long}|=45K$, and a smaller memory constraint with $|M|=10K$, $|M_{short}|=1K$, $|M_{long}|=9K$. For context, there are a total of 18.2M frames in the SAYCam training set and 2.5M frame in the KrishnaCam training set. Therefore, even the large memory constraint of 50K frames only stores 0.27% and 2.01% of the total training frames in the memory buffer for SAYCam and KrishnaCam respectively.
@@
  • mini-ImageNet classification. Following a similar evaluation protocol as Zhuang et al., [75], we evaluate the learned representations on a downstream classification task on a subsampled ImageNet [13] dataset (mini-ImageNet). We extract the features of the model and train a support vector machine (SVM) or a k-nearest neighbor (kNN) classifier to measure its classification performance. The mini-ImageNet dataset contains 20K training images and 5K test images across 100 classes.
  • ImageNet-1K classification. Similar to the evaluation protocol used in Purushwalkam et al., [43], we further evaluate the classification performance with a linear classifier on the larger ImageNet-1K [13] dataset with 1.28M training images and 50K test images across 1K classes.
  • Labeled-S classification. For SAYCam models, we evaluate the classification performance on the Labeled-S dataset [40]. The the Labeled-S dataset is a labeled subset of the SAYCam frames, containing a total of 5786 images across 26 classes after 10x subsampling of frames. We randomly use 50% as training data and 50% as test data.
  • OAK object detection. For KrishnaCam models, we evaluate the object detection performance on the Objects Around Krishna (OAK) dataset [60], which includes bounding box annotations of 105 object categories on a subset of the KrishnaCam frames. We fine-tune the model on the entire training set of OAK for 10 epochs before evaluating on the OAK validation set.
@@
We compare a number of competitive SSL methods for image and video representation learning, and different memory buffer strategies:
@@
  • SimCLR. In prior studies, Zhuang et al., [75] showed that SimCLR [11] is the strongest self-supervised learning method under streaming video setting, outperforming other SSL methods such as BYOL [19] and Barlow Twins [71].
@@
  • SimSiam. In prior work, Purushwalkam et al., [43] showed that SimSiam [12] is able to learn good representations from egocentric video data.
@@
  • Osiris. Osiris [73] is a state-of-the-art unsupervised continual learning method that is developed towards static image sequences.
@@
  • TC. Temporal classification (TC) [40] is a simple self-supervised learning method that is shown to work well on the SAYCam dataset under IID setting. It also uses temporal segments as a source of self-supervision; however, it does not actively group the frames together but instead relies on fixed intervals.
@@
  • Reservoir Sampling. We mainly use reservoir sampling [58] as a default baseline approach for updating the memory buffer, which uniformly samples from all the seen images in the memory.
@@
  • MinRed Buffer. The minimum redundancy (MinRed) buffer [43], also designed for the streaming setting, alleviates the temporal correlation of data in the continuous video stream by maintaining minimally redundant samples in the replay buffer.
@@
  • Two-tier Buffer. As in MemStoryboard, we use a long-term memory updated with reservoir sampling and short-term memory updated with first-in-first-out (FIFO), but we do not apply the temporal contrastive loss or the temporal segmentation module.
@@ 5.2 Main Results
In Tables 1 and 2, we report the main results on streaming SSL on both SAYCam and KrishnaCam. Firstly, we observe that all SSL methods work poorly in the streaming setting without replay, and larger memory leads to better performance. In terms of memory buffer strategies, our two-tier memory hierarchy and MinRed [43] outperforms reservoir sampling.
Memory Storyboard achieves superior performance in all readout tasks compared to other streaming SSL models. For SimCLR-based methods, Memory Storyboard outperforms the baseline Reservoir sampling method by an average of 1.85% on the SVM readout performance of mini-ImageNet classification and considerably narrows the gap between streaming learning and IID training. Memory Storyboard also significantly outperforms all baseline methods with a considerable gap by around 3% on AP50 on the challenging OAK object detection benchmark. For SimSiam-based methods, Memory Storyboard not only outperforms all streaming learning baselines by a considerable margin but also beats IID SimSiam training on all readout tasks when using a 50K replay buffer size.
Memory Storyboard with SimSiam achieves the overall best performance across different training datasets and evaluation metrics. We hypothesize that Memory Storyboard works better with SimSiam [12] than SimCLR [11] in our experiments due to the fact that SimCLR treats some highly correlated images in the same batch as negative samples during training, which hinders effective representation learning. This issue is exacerbated in the SAYCam experiments due to the high frequency (25 fps) of the SAYCam video stream. By incorporating the temporal contrastive loss in Memory Storyboard, we successfully address this issue by utilizing only images in other temporal classes as negative samples.
@@
We visualize the temporal segments produced Memory Storyboard at the end of training in Figure 3. The results demonstrate that the our temporal segmentation module can produce semantically meaningful temporal segments, showing its strong temporal abstraction capability. We emphasize that the representations are entirely developed during the streaming SSL training as the networks are trained from scratch.
We also visualize the object detection results produced by Memory Storyboard when fine-tuned on the OAK dataset [60] in Figure 4. We observe that the fine-tuned model can successfully detect objects in cluttered environments. The results show that the representations learned by Memory Storyboard can be effectively transferred to downstream tasks which requires more fine-grained features.
Figure 3: Visualization of the temporal segments produced by Memory Storyboard on (a) SAYCam (b)(c) KrishnaCam at the end of training. The images are sampled at 10 seconds per frame. Each color bar correspond to a temporal class (the first and the last class might be incomplete). Temporal segments produced at the beginning of training are provided in the appendix for comparison.
5.3 Other Training Factors
In this section, we study how varying different training factors affects the performance of Memory Storyboard, including subsampling rate, average segment length, and normalization layers. For evaluation on the downstream tasks, we use SVM readout top-1 accuracy for classification tasks and AP50 for fine-tuning on OAK. We use SimCLR [11] as the base SSL method for training.
@@ 5.3 Other Training Factors
We train Memory Storyboard with different subsampling rates when adding data fetched from the current stream to the short-term memory. Results are shown in Table 3. A subsampling ratio of 8 works best for SAYCam, while a ratio of 4 works best for KrishnaCam. Since the two datasets are decoded at different frequencies (25 fps for SAYCam and 10 fps for KrishnaCam), the effective frequency of frames entering the short-term buffer is 3.13 and 2.50 fps respectively. The result suggests that an effective frequency of around 3 fps is preferable although the optimal subsample ratio is dependent on nature of the video stream. Intuitively, when the subsampling ratio is too small, the images entering the short-term buffer may have too much temporal correlation and hence would hurt the performance; when the subsampling ratio is too big, the model skips too many frames without training on them and the temporal clustering may also become less precise.
@@
Ratio mini-ImageNet Labeled-S mini-ImageNet OAK AP50
1$\times$ 36.70 55.29 35.54 38.55
2$\times$ 37.18 55.43 35.60 37.38
4$\times$ 38.38 55.84 36.36 38.67
8$\times$ 39.58 56.29 35.48 38.90
16$\times$ 38.62 55.81 35.88 38.22
@@
Average Segment Length.
We trained Memory Storyboard with different average segment length $T$ ranging from 1 minute to 10 minutes on SAYCam and KrishnaCam. The results are shown in Table 4. We demonstrate that the performance of Memory Storyboard is generally robust to average segment length (which determines the number of temporal segments in the segmentation module). We also find that the performance on downstream tasks becomes worse when the average segment length is very long $(T=10\text{ min})$ on both datasets. This observation is different from that of temporal classification [40] which claims longer segments are more helpful.
@@
mini-ImageNet Labeled-S mini-ImageNet OAK AP50
@@
BatchNorm vs. GroupNorm.

### v3

@@ 5.1 Experiment Setup
We use two real-world egocentric video datasets in the experiments: (1) the child S subset of SAYCam dataset (Sullivan et al., 2021), which contains 221 hours of video data collected from a head-mounted camera on the child from age 6-32 months, decoded at 25 fps; (2) the KrishnaCam dataset (Singh et al., 2016), which contains 70 hours of video data spanning nine months of the life of a graduate student, decoded at 10 fps. These two datasets have also been adopted in a number of existing self-supervised learning literature (Orhan et al., 2020; Purushwalkam et al., 2022; Zhuang et al., 2022; Vong et al., 2024).
@@
Following the architectural choices of Osiris (Zhang et al., 2024), we use ResNet-50 (He et al., 2016) as the feature extractor with group normalization (Wu & He, 2018) and the Mish activation function (Misra, 2020). Unless otherwise specified, the default hyperparameter values we use in our experiments are $b=64$, $B=512$, $T=4.5K$ for SAYCam and $T=1.8K$ for KrishnaCam (both corresponding to 3 minutes of raw video), subsampling rate $r=8$ for SAYCam and $r=4$ for KrishnaCam. We train the models with two sets of memory sizes to evaluate their performance across different memory constraints: a larger memory constraint with $|M|=50K$, $|M_{short}|=5K$, $|M_{long}|=45K$, and a smaller memory constraint with $|M|=10K$, $|M_{short}|=1K$, $|M_{long}|=9K$. For context, there are a total of 18.2M frames in the SAYCam training set and 2.5M frames in the KrishnaCam training set. Therefore, even the large memory constraint of 50K frames only stores 0.27% and 2.01% of the total training frames in the memory buffer for SAYCam and KrishnaCam respectively. For the main experiments (Tables 1 and 2), we employ label merging with $\tau=0.998$ (i.e., the similarity threshold $\delta$ is the top $0.002$ quantile of the off-diagonal values in the similarity matrix); for all other experiments, label merging is not employed unless explicitly specified otherwise. Each experiment is run on one A100 GPU.
@@
  • mini-ImageNet Classification: Following a similar evaluation protocol as Zhuang et al. (2022), we evaluate the learned representations on downstream classification of a subsampled ImageNet (Deng et al., 2009) dataset (mini-INet). We extract the features of the model and train a support vector machine (SVM) to measure its classification performance. The mini-ImageNet dataset contains 20K training images and 5K test images across 100 classes.
  • ImageNet-1K and iNaturalist Classification: Similar to the evaluation protocol used in Purushwalkam et al. (2022), we further evaluate the classification performance with a linear classifier on the larger ImageNet-1K (Deng et al., 2009) dataset (INet) with 1.28M training images and 50K test images across 1K classes, and the iNaturalist-2018 (Van Horn et al., 2018) dataset (iNat) with 437K training images and 24K test images across 8142 classes.
  • Labeled-S Classification: For SAYCam models, we evaluate the classification performance on the Labeled-S dataset, following Orhan et al. (2020). The Labeled-S dataset is a labeled subset of the SAYCam frames, containing a total of 5786 images across 26 classes after 10x subsampling of frames. We randomly use 50% as training data and 50% as test data.
  • OAK Object Detection: For KrishnaCam models, we evaluate the object detection performance on the Objects Around Krishna (OAK) dataset (Wang et al., 2021), which includes bounding box annotations of 105 object categories on a subset of the KrishnaCam frames. We fine-tune the model on the entire training set of OAK for 10 epochs before evaluating on the OAK validation set, and report the AP50 metric.
@@
We compare Memory Storyboard with a number of competitive SSL methods for image and video representation learning, and different memory buffer strategies:
@@
  • SimCLR: In prior studies, Zhuang et al. (2022) showed that SimCLR (Chen et al., 2020) is the strongest self-supervised learning method under streaming video setting, outperforming other SSL methods such as BYOL (Grill et al., 2020) and Barlow Twins (Zbontar et al., 2021).
@@
  • SimSiam: In prior work, Purushwalkam et al. (2022) showed that SimSiam (Chen & He, 2021) is able to learn good representations from egocentric video data.
@@
  • Osiris: Osiris (Zhang et al., 2024) is a state-of-the-art unsupervised continual learning method that is developed towards static image sequences.
@@
  • TC: Temporal classification (TC) (Orhan et al., 2020) is a simple self-supervised learning method that is shown to work well on the SAYCam dataset under IID setting. It also uses temporal segments as a source of self-supervision; however, it does not actively group the frames together but instead relies on fixed intervals.
@@
  • Reservoir Sampling: We mainly use reservoir sampling (Vitter, 1985) as a default baseline approach for updating the memory buffer, which uniformly samples from all the seen images in the memory.
@@
  • MinRed Buffer: The minimum redundancy (MinRed) buffer (Purushwalkam et al., 2022) is a streaming self-supervised learning algorithm that alleviates temporal correlations in continuous video streams by storing minimally redundant samples in the replay buffer.
@@
  • Two-tier Buffer: As in MemStoryboard, we use a long-term memory updated with reservoir sampling and short-term memory updated with first-in-first-out (FIFO), but we do not apply the temporal contrastive loss or the temporal segmentation module.
Method mini-INet INet iNat Labeled-S mini-INet INet iNat Labeled-S
IID SimCLR (Chen et al., 2020) 44.04 30.44 8.69 59.50 44.04 30.44 8.69 59.50
IID SimSiam (Chen & He, 2021) 29.02 20.92 4.91 42.71 29.02 20.92 4.91 42.71
SimCLR No Replay 5.76 2.22 0.07 19.13 5.76 2.22 0.07 19.13
SimSiam No Replay 6.44 1.47 0.04 22.03 6.44 1.47 0.04 22.03
Replay - 10k Replay - 50k
Osiris (Zhang et al., 2024) 31.16 19.48 4.68 45.81 36.90 23.16 5.85 50.88
TC (Orhan et al., 2020) 33.92 19.03 5.84 48.09 36.68 22.72 8.24 52.22
SimCLR (Chen et al., 2020) 33.02 20.13 4.74 49.29 37.96 23.75 6.91 53.67
   +MinRed (Purushwalkam et al., 2022) 33.62 20.21 5.12 48.88 38.66 24.10 6.81 54.75
   +Two-tier (Ours) 33.80 20.70 5.61 49.05 39.22 24.93 7.07 55.43
   +MemStoryboard (Ours) 34.18 22.59 6.34 51.09 38.84 26.87 8.17 56.26
SimSiam (Chen & He, 2021) 20.90 13.72 2.55 39.12 26.66 14.44 3.79 43.09
   +MinRed (Purushwalkam et al., 2022) 22.68 17.85 3.17 39.78 25.58 18.99 4.24 40.37
   +Two-tier (Ours) 21.78 16.87 2.76 39.19 28.34 20.24 3.99 42.95
   +MemStoryboard (Ours) 36.86 26.70 8.46 49.87 41.46 28.92 10.41 53.78
Table 1: Results on streaming SSL from SAYCam (Sullivan et al., 2021). Downstream evaluation on object classification (Accuracy %) for SSL models trained under the streaming setting. For “No Replay” and “IID” the results are the same for different memory buffer sizes. The “IID” methods are not under the streaming setting and are for reference only as a performance “upper bound” with the same number of gradient updates. Unless specified, standard reservoir sampling is used in the replay buffer.
Method mini-INet INet iNat OAK mini-INet INet iNat OAK
IID SimCLR (Chen et al., 2020) 36.90 23.77 5.60 39.54 36.90 23.77 5.60 39.54
IID SimSiam (Chen & He, 2021) 28.58 22.28 4.16 44.86 28.58 22.28 4.16 44.86
SimCLR No Replay 4.84 1.35 0.07 14.01 4.84 1.35 0.07 14.01
SimSiam No Replay 8.88 1.92 0.05 27.34 8.88 1.92 0.05 27.34
Replay - 10k Replay - 50k
Osiris (Zhang et al., 2024) 30.10 19.03 3.55 32.25 32.38 20.85 3.96 33.78
TC (Orhan et al., 2020) 32.58 19.19 6.01 32.61 32.94 20.50 6.25 28.56
SimCLR (Chen et al., 2020) 31.46 19.09 4.43 31.92 34.98 22.37 5.19 33.30
   +MinRed (Purushwalkam et al., 2022) 31.56 19.93 4.69 34.78 34.84 22.29 5.30 35.65
   +Two-tier (Ours) 33.26 20.39 5.04 33.72 35.78 22.42 5.29 35.68
   +MemStoryboard (Ours) 33.30 22.52 5.65 36.01 36.08 25.37 6.28 38.58
SimSiam (Chen & He, 2021) 19.16 12.94 2.85 39.38 21.84 14.13 3.56 41.13
   +MinRed (Purushwalkam et al., 2022) 20.90 14.53 3.16 43.74 22.88 17.64 5.12 44.17
   +Two-tier (Ours) 20.08 13.76 2.91 43.68 22.14 17.06 3.73 44.41
   +MemStoryboard (Ours) 33.22 23.76 6.52 45.18 34.62 25.52 6.78 46.17
Table 2: Results on streaming SSL from KrishnaCam (Singh et al., 2016). Downstream evaluation on object classification (Accuracy %) and object detection (AP50 %) for SSL models trained under the streaming setting. The structure of the table is otherwise similar to Table 1.
@@ 5.2 Main Results
In Tables 1 and 2, we report the main results on streaming SSL on both SAYCam and KrishnaCam. Firstly, we observe that all SSL methods work poorly in the streaming setting without replay, and larger memory leads to better performance. In terms of memory buffer strategies, our two-tier memory hierarchy and MinRed (Purushwalkam et al., 2022) outperform reservoir sampling.
Memory Storyboard achieves superior performance in all readout tasks compared to other streaming SSL models. For SimCLR-based methods, Memory Storyboard considerably narrows the gap between streaming learning and IID training. Memory Storyboard also significantly outperforms all baseline methods with a considerable gap on both the ImageNet classification and the challenging OAK object detection benchmark. For SimSiam-based methods, Memory Storyboard not only outperforms all streaming learning baselines by a considerable margin but also beats IID SimSiam training on all readout tasks.
Memory Storyboard with SimSiam achieves the overall best performance across different training datasets and evaluation metrics. We hypothesize that Memory Storyboard works better with SimSiam (Chen & He, 2021) than SimCLR (Chen et al., 2020) in our experiments because the SimCLR loss can be conflicting with the temporal contrastive objective. SimCLR treats some highly correlated images in the same batch as negative samples during training, despite the fact that the other images in the batch sampled from the short-term buffer might be very similar to the current image since they are temporally close to each other. This issue is exacerbated in the SAYCam experiments due to the high frequency (25 fps) of the SAYCam video stream. By incorporating the temporal contrastive loss in Memory Storyboard, we successfully address this issue by utilizing only images in other temporal classes as negative samples.
@@
We visualize the temporal segments produced by Memory Storyboard at the end of training in Figure 3. The results demonstrate that our temporal segmentation module can produce semantically meaningful temporal segments, showing its strong temporal abstraction capability. We emphasize that the representations are entirely developed during the streaming SSL training as the networks are trained from scratch. We provide additional qualitative results in Appendix C.
Figure 3: Visualization of the temporal segments produced by Memory Storyboard on (a) SAYCam (b)(c) KrishnaCam at the end of training. The images are sampled at 10 seconds per frame. Each color bar corresponds to a temporal class (the first and the last class might be incomplete). Temporal segments produced at the beginning of training are provided in the appendix for comparison.
5.3 Ablation Experiments and Other Training Factors
In this section, we study how varying different training factors affect the performance of Memory Storyboard, including label merging, subsampling rate, and average segment length. We use SimCLR (Chen et al., 2020) as the base SSL method for training, and a long-term memory size of 50K unless otherwise specified. Please refer to Appendix C for additional ablation experiments.
Base SSL Method Label Merging mini-INet INet iNat Labeled-S mini-INet INet iNat Labeled-S
Replay - 10k Replay - 50k
SimCLR 35.02 20.72 5.65 51.33 39.58 24.78 7.77 56.29
SimCLR 34.18 22.59 6.34 51.09 38.84 26.87 8.17 56.26
SimSiam 36.72 22.99 6.66 49.12 41.32 26.37 9.85 53.29
SimSiam 36.86 26.70 8.46 49.87 41.46 28.92 10.41 53.78
Table 3: Ablation on label merging for MemStoryboard trained on SAYCam.
Base SSL Method Label Merging mini-INet INet iNat OAK mini-INet INet iNat OAK
Replay - 10k Replay - 50k
SimCLR 33.72 20.13 5.64 35.77 36.36 22.75 6.10 38.67
SimCLR 33.30 22.52 5.65 36.01 36.08 25.37 6.28 38.58
SimSiam 33.78 21.38 6.51 45.33 35.20 22.75 6.71 46.64
SimSiam 33.22 23.76 6.52 45.18 34.62 25.52 6.78 46.17
Table 4: Ablation on label merging for MemStoryboard trained on KrishnaCam.
Label Merging.
We train MemStoryboard with and without the label merging mechanism and present the results in Tables 3 and 4. We observe that incorporating label merging consistently improves performance on the ImageNet and iNaturalist classification tasks, which are out-of-domain classification settings. This suggests that grouping labels for representation learning is helpful in recognizing new classes at test time. While the improvements are generally modest in absolute terms, they are robust across different datasets and buffer sizes. Performance on the other benchmarks (mini-ImageNet, Labeled-S, and OAK) remains largely stable with or without label merging.
@@ 5.3 Ablation Experiments and Other Training Factors
We train Memory Storyboard with different subsampling rates when adding data fetched from the current stream to the short-term memory. Results are shown in Table 5. A subsampling ratio of 8 works best for SAYCam, while a ratio of 4 works best for KrishnaCam. Since the two datasets are decoded at different frequencies (25 fps for SAYCam and 10 fps for KrishnaCam), the effective frequency of frames entering the short-term buffer is 3.13 and 2.50 fps respectively. The result suggests that an effective frequency of around 3 fps is preferable although the optimal subsample ratio is dependent on the nature of the video stream. Intuitively, when the subsampling ratio is too small, the images entering the short-term buffer may have too much temporal correlation and hence would hurt the performance; when the subsampling ratio is too big, the model skips too many frames without training on them and the temporal clustering may also become less precise.
@@
Ratio mini-INet Labeled-S mini-INet OAK AP50
[…]
@@
mini-INet Labeled-S mini-INet OAK AP50
@@
Average Segment Length.
We trained Memory Storyboard with different average segment lengths $T$ ranging from 1 minute to 10 minutes on SAYCam and KrishnaCam. The results are shown in Table 6. We demonstrate that the performance of Memory Storyboard is generally robust to average segment length (which determines the number of temporal segments in the segmentation module). We also find that the performance on downstream tasks becomes worse when the average segment length is very long $(T=10\text{ min})$ on both datasets. This observation is different from that of temporal classification (Orhan et al., 2020) which claims longer segments are more helpful.
Figure 4: Memory Storyboard model performance on SAYCam with different long-term memory sizes (5k, 10k, 50k, and 100k) and varying training batch compositions (12.5% – 75.0% from $M_{short}$) using SVM readout. Each colored line represents the performance of different training batch compositions when the model has seen the same amount of data from the stream. Each black line represents the performance of different training batch compositions when the model has taken the same number of gradient updates.
5.4 Optimal Batch Composition Under Different Memory Constraints
In Memory Storyboard, the training batch is composed of samples from both the long-term memory and the short-term memory (see Figure 2). However, the optimal composition ratio of the training batch, i.e. the optimal percentage of data in the training batch that comes from the short-term memory, is yet to be explored. Sampling more data from the short-term memory means we can digest more data within a fixed number of training steps, but there will be more distribution shift between different training batches. On the other hand, sampling more data from the long-term memory buffer may result in overfitting on the long-term memory data. In this section, we experiment with different memory sizes and training composition and demonstrate the optimal batch composition under different memory constraints.
We fix the size of the short-term memory $|M_{short}|$ to be $5K$ and vary the memory constraint for the long-term memory $|M_{long}|=5K,10K,50K,100K$. For each long-term memory size, we experiment with batch size from data stream $b=64,128,192,256,320,384$ (which corresponds to 12.5% though 75% of the training batch size). We sample $b$ images from the short-term memory and 512b512-b images from the long-term memory to compose a training batch. We evaluate the model with SVM readout on mini-ImageNet after the model has seen every 10% of the entire data stream and plot the results in Figure 4. We discuss the different observations for large memory size and small memory size respectively.
  • Large $|M_{long}|$: When $|M_{long}|$ is large (Figures 4(c) and 4(d)), overfitting on the memory is unlikely and hence we can sample more data from the long-term memory and the performance still keeps increasing as the model sees more data. Hence, with the same amount of data seen by the model (colored curves), it is better to sample only a small batch from the short-term memory. However, when we control the number of model update steps to the same (black curves), neither focusing on the short-term memory nor focusing on the long-term memory is preferable. In such cases, the optimal batch size from the short-term memory is at roughly 50% of the training batch.
  • Small $|M_{long}|$: When $|M_{long}|$ is small (Figures 4(a) and 4(b)), the model is prone to overfitting on the memory. As a result, with the same number of model update steps (black curves), taking more images from $M_{short}$ gives better results. With the same amount of data seen by the model (colored curves), getting a higher percentage of data from $M_{long}$ has an advantage in the beginning when there is less memory overfitting. Ultimately, focusing on $M_{short}$ is more beneficial in the late stage.
To summarize, the optimal training batch composition depends on memory and compute constraints. More samples from the long-term memory are preferred when a large memory is available (e.g., 50K images from a 200-hour stream) and model performance is evaluated after seeing a fixed amount of data. More samples from the short-term memory are preferred when memory is limited to prevent overfitting on the long-term memory data. When both memory and compute are sufficient and performance is measured under a fixed compute budget, a balanced batch composition is most effective for real-time learning.
6 Conclusion
The ability to continuously learn from large-scale uncurated streaming video data is crucial for applying self-supervised learning methods in real-world embodied agents. Existing works have limited exploration of this problem, have mainly focused on static datasets, and do not perform well in the streaming video setting. Inspired by the event segmentation mechanism in human cognition, in this work, we propose Memory Storyboard, which leverages temporal segmentation to produce a two-tier memory hierarchy akin to the short-term and long-term memory of humans. Memory Storyboard combines a temporal contrastive objective and a standard self-supervised contrastive objective to facilitate representation learning from scratch through streaming video experiences. Memory Storyboard achieves state-of-the-art performance on downstream classification and object detection tasks when trained on real-world large egocentric video datasets. By studying the effects of subsampling rates, average segment length, normalization, and optimal batch composition under different compute and memory constraints, we also offer valuable insights on the design choices for streaming self-supervised learning.
Acknowledgements
We thank James McClelland for a helpful discussion on event segmentation and pointers to relevant literature. We thank Peiqi Liu and Anh Ta for their explorations on training self-supervised learning models on the SAYCam dataset. The work is supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) under grant RS-2024-00469482, funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. YY is supported by the Meta AI Mentorship Program. The compute is supported through the NYU IT High Performance Computing resources, services, and staff expertise.
Appendix A Experiment Details
Model Architecture.
On top of the ResNet backbone, we use a two-layer MLP with 2048 hidden units, 128 output units, and ReLU activation function as the projector. In Memory Storyboard, we create two separate projectors for $\mathcal{L}_{TCL}$ and $\mathcal{L}_{SSL}$.
Training.
For all experiments in Tables 1 and 2, we used a total batch size of 512 (64 from $M_{short}$ and 448 from $M_{long}$ by default). The input resolution of the images to the model is 112. We apply a standard data augmentation pipeline for SSL methods following Zhuang et al. (2022), which include random resized crop, random horizontal flip, random color jitter, random grayscale, random Gaussian filter, and color-normalization with ImageNet (Deng et al., 2009). For the SimCLR (Chen et al., 2020), Osiris (Zhang et al., 2024), and TC (Orhan et al., 2020) experiments, we used the Adam (Kingma & Ba, 2015) optimizer with a constant learning rate of 0.001, and a projector with 2 MLP layers of size 2048 and 128 respectively. For the SimSiam (Chen & He, 2021) experiments, we used the SGD optimizer with learning rate 0.05, momentum 0.9, and weight decay 1e-4, and a projector with 3 MLP layers of size 2048.
Evaluation.
For mini-ImageNet and Labeled-S evaluations, the streaming SSL models are evaluated every 5% of the entire dataset. That is, we store 20 model checkpoints throughout the streaming training and evaluate them on mini-ImageNet and Labeled-S with SVM readout. The best results among these checkpoints are reported. Similar to Zhuang et al. (2022), for SVM readout, we report the best performance among learning rate values {1e-7, 1e-6, 1e-5, 1e-4, 1e-3, 1e-2, 1e-1, 1, 1e1, 1e2}.
For ImageNet-1K and iNaturalist-2018 evaluations, we evaluate the final model after streaming SSL training on the entire dataset. Following Purushwalkam et al. (2022), we train a linear classifier on top of the normalized learned representations and report the classification accuracy. We used a batch size of 1024. For ImageNet-1K, we used the LARS (You et al., 2017) optimizer with learning rate 3.0, momentum 0.9, and cosine learning rate schedule for 10 epochs. For iNaturalist-2018, we used the LARS (You et al., 2017) optimizer with learning rate 12.0, momentum 0.9, and cosine learning rate schedule for 20 epochs.
For OAK evaluations, we use Faster R-CNN (Ren et al., 2015), a popular two-stage object detector. We initialize the ResNet-50 (He et al., 2016) backbone with the backbone of the final checkpoint of the streaming SSL model, and fine-tune the entire model on OAK with IID training for 10 epochs, following the training configurations of (Wu et al., 2023).
Appendix B Additional Related Work
Self-Supervised Learning.
A large number of self-supervised representation learning methods in computer vision follows the contrastive learning framework (Oord et al., 2018; Misra & Maaten, 2020; Tian et al., 2020; He et al., 2020; Chen et al., 2020; Chen & He, 2021) which maximizes the agreement of representations of two augmented views of the same image and minimizes that of different images. Extending this idea, the supervised contrastive (SupCon) method (Khosla et al., 2020) uses the labels as an extra supervision signal to get multiple positive crops for each anchor image. Other recent self-supervised learning works include pretext tasks (Doersch et al., 2015; Noroozi & Favaro, 2016; Gidaris et al., 2018; Pathak et al., 2016), feature space clustering (Caron et al., 2018; 2020; Ren et al., 2021), distillation with asymmetric architectures (Grill et al., 2020; Chen & He, 2021), redundancy reduction (Zbontar et al., 2021; Bardes et al., 2022), and masked autoencoding (He et al., 2022). Most relevant of these to our work, Orhan et al. (2020) proposes the temporal classification objective, which outperforms contrastive learning objectives on the SAYCam dataset (Sullivan et al., 2021). Our work enhances the temporal classification method by using a more flexible supervised contrastive objective, and leveraging temporal segmentation (Potapov et al., 2014; Afham et al., 2023), which have been used extensively in video summarization (Zhu et al., 2020; Zhang et al., 2016; Rochan et al., 2018).
Temporal Segmentation in Human Cognition.
Prior research in psychology and cognitive sciences has shown that humans, including infants, are able to identify boundaries between action segments (Newtson et al., 1977; Zacks et al., 2001; Baldwin et al., 2001; Saylor et al., 2007; Baldassano et al., 2017; Yates et al., 2022). Evidence in neuro-imaging further shows that event segmentation is an automatic component in human perception (Zacks & Swallow, 2007). Temporal event segmentation has proven to be critical for memory formation and retrieval (Lassiter & Slaw, 1991; Ezzyat & Davachi, 2011; DuBrow & Davachi, 2013; Silva et al., 2019; Sasmita & Swallow, 2022). The temporal segmentation component in our proposed framework is motivated by how humans interpret videos as segments with coherent semantics. We demonstrate that temporal segmentation can improve the learned visual representation.
Appendix C Additional Results
C.1 Performance of Different Normalization Layers